# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ 5.1 Robust estimation with treatment overlap but without recommendation overlap
We next describe methods for robust extrapolation under structural assumptions about the smoothness of the outcome models. Under violations of overlap, the only unknown quantity is $p_{t\mid r}(X)$ in regions of no overlap in recommendation; however, a plausible assumption is that the underlying function is smooth in covariates. A robust approach obtains worst-case bounds on policy value under all functions compatible with a particular smoothness assumption. On the other hand, we assume that overlap holds with respect to $T$ given covariates $X$, so our finer-grained approach via Assumption˜2 yields milder penalties due to robustness since we need robustly extrapolate only the treatment response to recommendations, $p_{t\mid r}(X),$ rather than the outcome model’s $\mu_{t}(X).$ Define the regions of no overlap as the following: Let $\mathcal{X}^{\text{nov}}_{r}=\{x:P(R=r\mid x)=0\}$; in this region, we do not jointly observe all potential values of $(t,r,x)$. In addition, let $\mathcal{X}^{\text{nov}}=\bigcup_{r}\mathcal{X}^{\text{nov}}_{r}$. Correspondingly, define the overlap region as $\mathcal{X}^{\text{ov}}=(\mathcal{X}^{\text{nov}})^{c}.$ We consider uncertainty sets for ambiguous treatment recommendation probabilities. For example, one plausible structural assumption is monotonicity of treatment in recommendation. We define the following uncertainty set:
$\mathcal{U}_{q_{t\mid r}}\coloneqq\left\{q_{1\mid r}(x^{\prime})\colon q_{1\mid r}(x)\geq p_{1\mid r}(x),\;\forall x\in\mathcal{X}^{\text{nov}}_{r}\;\sum_{t\in\mathcal{T}}q_{t\mid r}(x)=1,\forall x,r\right\}$
We could assume uniform bounds on unknown probabilities; more refined bounds, such as Lipschitz smoothness with respect to some distance metric $d$; or boundedness.
$\textstyle\mathcal{U}_{\text{lip}}$ $\coloneqq\left\{q_{1\mid r}(x^{\prime})\colon d(q_{1\mid r}(x^{\prime}),p_{1\mid r}(x))\leq Ld(x^{\prime},x),\;(x^{\prime},x)\in(\mathcal{X}^{\text{nov}}\times\mathcal{X}^{\text{nov}})\right\}$
@@
Define $V_{ov}(\pi)\coloneqq\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)p_{t\mid r}(X)\mu_{t}(X)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}]$. Let $\mathcal{U}$ denote the uncertainty set including any custom constraints, e.g., $\mathcal{U}=\mathcal{U}_{q_{t\mid r}}\cap\mathcal{U}_{\text{lip}}$. Then, we may obtain robust bounds by optimizing over regions of no overlap:
$\textstyle\overline{V}(\pi)$ $\coloneqq V_{ov}(\pi)+\overline{V}_{nov}(\pi),\;\;$
@@
$\text{ where }\overline{V}_{nov}(\pi)$ $\textstyle\coloneqq\underset{q_{tr}(X)\in\mathcal{U}}{\max}\left\{\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)\mu_{t}(X)q_{tr}(X)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]]\right\}$
Lemma 1 (Binary outcomes, constant bound).
Let $\mathcal{U}_{\text{cbnd}}\coloneqq\left\{q_{t\mid r}(x^{\prime})\colon\underline{B}\leq q_{1\mid r}(x^{\prime})\leq\overline{B}\right\}$ and $\mathcal{U}=\mathcal{U}_{q_{t\mid r}}\cap\mathcal{U}_{\text{cbnd}}.$ Define $\beta_{t\mid r}\coloneqq\mathbb{E}[q_{t\mid r}(X,A)\mid T=t].$If $T\in\{0,1\},$
$\overline{V}_{no}(\pi)$ $=\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[c^{*}_{rt}\beta_{t\mid r}\mathbb{E}[Y\pi_{r}(X)\mid T=t]\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]],$
@@
We consider the case of continuous-valued outcomes in the example setting of the simple resource parity–constrained program of LABEL:eqn-resource-parity-constrained. We first study simple uncertainty sets, such as intervals, to deduce insights about the robust policy, with a more general reformulation in the appendix.
Proposition 6 (Robust linear program).
Suppose that $r,t\in\{0,1\}$ and $q_{r1}(\cdot,u)\in\mathcal{U}_{bnd},\forall r,u.$ Define
$\tau(x,a)$ $\coloneqq\mu_{1}(x,a)-\mu_{0}(x,a),\;\;\Delta B_{r}(x,u)\coloneqq(\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)),$
$B^{\text{mid}}_{r}(x,u)$ $\coloneqq\underline{B}_{r}(x,u)+\frac{1}{2}\Delta B_{r}(x,u),\;\;c_{1}(\pi)\textstyle\coloneqq\sum_{r}\mathbb{E}[\tau\pi_{r}B^{\text{mid}}],$
$\mathbb{E}[\Delta_{ov}T(\pi)]$ $\coloneqq\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}\{X\in\mathcal{X}^{\text{ov}}\}\mid A=b]\;\;$
Then, the robust linear program is:
$\min\;$ $\textstyle V_{ov}(\pi)+\mathbb{E}[\mu_{0}]+c_{1}(\pi)-\frac{1}{2}\sum_{r}\mathbb{E}[\left|\tau\right|\pi_{r}\Delta B_{r}(X,A)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}]$
s.t. $\textstyle\sum_{r}\{\mathbb{E}[\pi_{r}\overline{B}_{r}(X,A)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=a]-\mathbb{E}[\pi_{r}\underline{B}_{r}(X,A)\mathbb{I}\{X\in\mathcal{X}^{\text{nov}}\}\mid A=b]\}+\Delta_{ov}^{T}(\pi)\leq\epsilon$
6 Additional Fairness Constraints and Policy Optimization
We previously discussed policy optimization over unrestricted decision rules given estimates. We now introduce a general methodology to handle 1) optimization over a policy class of restricted functional form and 2) more general fairness constraints. We first introduce the fair-classification algorithm of Agarwal et al. (2018) and then describe our extensions to obtain variance-sensitive regret bounds and less conservative policy optimization (using a regularized empirical risk minimization (ERM) argument given in Chernozhukov et al., 2019).
Algorithm and setup
Algorithm 1 $\text{REDFAIR}(\mathcal{D},g,\mathcal{E},M,d)$
1: Input: $\mathcal{D}=\left\{\left(X_{i},R_{i},T_{i},Y_{i},A_{i}\right)\right\}_{i=1}^{n}$, $g,\mathcal{E},{M},\widehat{{d}}$, B, accuracy $\nu$, $\alpha$, stepsize $\omega,$ initialization ${\theta}_{1}={0}\in\mathbb{R}^{|\mathcal{K}|}$
2:for iteration $t=1,2,\dots$ do
3:  Set $\lambda_{t,k}=B\frac{\exp\left\{\theta_{k}\right\}}{1+\sum_{k^{\prime}\in\mathcal{X}}\exp\left\{\theta_{k^{\prime}}\right\}}$ for all $k\in\mathcal{K}$, $\beta_{t}\leftarrow\operatorname{BEST}_{\beta}\left({\lambda}_{t}\right)$, $\widehat{Q}_{t}\leftarrow\frac{1}{t}\sum_{t^{\prime}=1}^{t}\beta_{t^{\prime}}$ $\hat{{\lambda}}_{t}\leftarrow\frac{1}{t}\sum_{t^{\prime}=1}^{t}{\lambda}_{t^{\prime}},$
4:  $\bar{L}\leftarrow L(\hat{Q}_{t},\operatorname{BEST}_{{\lambda}}(\hat{Q}_{t})),\underline{L}\leftarrow(\operatorname{BEST}_{\beta}(\hat{{\lambda}}_{t}),\hat{{\lambda}}_{t})$,
5:  $\nu_{t}\leftarrow\max\{L(\hat{Q}_{t},\hat{{\lambda}}_{t})-\underline{L},\quad\bar{L}-L(\hat{Q}_{t},\hat{{\lambda}}_{t})\}$, If $\nu_{t}\leq\nu$ then return $(\hat{Q}_{t},\hat{{\lambda}}_{t})$
6:  $\theta_{t+1,i}=\theta_{t}+\log(1-\omega({M}\hat{{\mu}}\left(h_{t}\right)-\hat{{c}})_{j}),\forall i$
7:end for
We first describe the reductions-based approach for fair classification of Agarwal et al. (2018) before describing our adaptation for constrained policy learning and localized two-stage variance reduction. They consider classification (i.e., loss minimization) under fairness constraints that can be represented generically as a linear program. In the following, note that, to be consistent with standard form for linear programs, we consider costs $Y$ so that we can phrase the saddle point as minimization-maximization. The $|\mathcal{K}|$ linear constraints and $J$ groups (values of protected attribute $A$) are summarized via a coefficient matrix ${M}\in\mathbb{R}^{K\times J}$, which multiplies a vector of constraint moments $h_{j}(\pi),j\in[J]$ (with $J$ being the number of groups); $O=(X,A,R,T,Y)$ denotes our data observations and $d$ the constraint constant vector:
$h_{j}(\pi)=\mathbb{E}\left[g_{j}(O,\pi(X))\mid\mathcal{E}_{j}\right]\;\;\text{ for }j\in J,\;\;{M}{h}(\pi)\leq{d}$
The elements of $h_{j}(\pi)$ are average functionals (for example, the average treatment takeup in group $j$). Importantly, the moment function $g_{j}$ depends on $\pi$, while the conditioning event $\mathcal{E}_{j}$ cannot depend on $\pi$. Many important fairness constraints can nonetheless be written in this framework, such as burden/resource parity and parity in true positive rates, but not measures such as calibration whose conditioning event does depend on $\pi$. (See Section˜C.2 for examples omitted for brevity.)
Our objective function is the policy value $V(\pi)$. (Later, this is linearized (as in Agarwal et al., 2018) by optimizing over distributions over policies.) We further consider a convexification of $\Pi$ via randomized policies $Q\in\Delta(\Pi)$, where $\Delta(\Pi)$ is the set of distributions over $\Pi,$ i.e., a randomized classifier that samples a policy $\pi\sim Q$. Therefore, our target estimand is the optimal distribution $Q$ over policies $\pi$ that minimizes the objective value $V(Q)$ subject to the fairness constraints encoded in ${M}{h}(Q)\leq{d}$:
$\min_{Q\in\Delta(\Pi)}\{V(Q)\colon\;\;{M}{h}(Q)\leq{d}\}$
Next, we discuss the cost-weighted classification reduction of off-policy learning (Zhao et al., 2012), which we use to solve constrained off-policy learning via Agarwal et al. (2014).
@@ 6 Additional Fairness Constraints and Policy Optimization
We use a well-known reduction of policy learning to cost-sensitive classification, described in Section˜C.2.1 of the appendix. Therefore, the centered regret can be reparametrized via the parameter $\beta$ as: $J(\beta)=J(\operatorname{sgn}(f_{\beta}(\cdot)))=\mathbb{E}[\operatorname{sgn}(f_{\beta}(X))\left\{\psi\right\}].$ We can apply the standard reduction to cost-sensitive classification since $\psi_{i}\operatorname{sgn}(f_{\beta}(X_{i}))=\left|\psi_{i}\right|(1-2\mathbb{I}\left[\operatorname{sgn}(f_{\beta}(X_{i}))\neq\operatorname{sgn}(\psi_{i})\right]$). Then, we can use surrogate losses for the zero-one loss. Although many functional forms for $\ell(\cdot)$ are Fisher consistent, one such choice of $\ell$ is the logistic (cross-entropy) loss $\mathbb{E}[\left|\psi\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi))],\qquad l(g,s)=2\log(1+\exp(g))-(s+1)$.
Algorithm 2 Two-stage localized fair classification via reductions
1: Randomly split the data into two folds, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
2: Obtain $\hat{Q}^{*}_{1}$ and the index set of binding constraints $\hat{\mathcal{I}}_{1}$ by learning nuisances $\eta_{1}$ and running Algorithm˜1 on $\mathcal{D}_{1}$ with $\text{REDFAIR}(\mathcal{D}_{1},h,{\mathcal{E}},M,d;\eta_{1}).$

### v5

@@ 5.1 Extension: Algorithmic advice with treatment, but not recommendation overlap
We next describe methods for robust extrapolation under structural assumptions about the smoothness of the outcome models. Under violations of overlap, the only unknown quantity is $p_{t\mid r}(X)$ in regions of no overlap in recommendation; however, a plausible assumption is that the underlying function is smooth in covariates. A robust approach obtains worst-case bounds on policy value under all functions compatible with a particular smoothness assumption. On the other hand, we assume that overlap holds with respect to $T$ given covariates $X$, so our finer-grained approach via Section 3 yields milder penalties due to robustness since we need robustly extrapolate only the treatment response to recommendations, $p_{t\mid r}(X),$ rather than the outcome model’s $\mu_{t}(X).$ Define the regions of no overlap as the following: Let $\mathcal{X}^{\text{nov}}_{r}=\{x:P(R=r\mid x)=0\}$; in this region, we do not jointly observe all potential values of $(t,r,x)$. In addition, let $\mathcal{X}^{\text{nov}}=\bigcup_{r}\mathcal{X}^{\text{nov}}_{r}$. Correspondingly, define the overlap region as $\mathcal{X}^{\text{ov}}=(\mathcal{X}^{\text{nov}})^{c}.$ We consider uncertainty sets for ambiguous treatment recommendation probabilities. For example, one plausible structural assumption is monotonicity of treatment in recommendation, or other global shape restrictions. We could assume uniform bounds on unknown probabilities; more refined bounds, such as Lipschitz smoothness with respect to some distance metric $d$; or boundedness.
$\textstyle\mathcal{U}_{\text{lip}}$ $\coloneqq\left\{q_{1\mid r}(x^{\prime})\colon d(q_{1\mid r}(x^{\prime}),p_{1\mid r}(x))\leq Ld(x^{\prime},x),\;(x^{\prime},x)\in(\mathcal{X}_{\mathrm{nov}}\times\mathcal{X}_{\mathrm{ov}})\right\}$
@@
We write $\mathbb{I}_{\mathrm{ov}}(X):=\mathbb{I}\{X\in\mathcal{X}_{\mathrm{ov}}\}$ and $\mathbb{I}_{\mathrm{nov}}(X):=\mathbb{I}\{X\in\mathcal{X}_{\mathrm{nov}}\}$ for the overlap and no-overlap indicators. Define $V_{ov}(\pi)\coloneqq\sum_{t\in\mathcal{T},r\in\{0,1\}}\mathbb{E}[\pi_{r}(X)p_{t\mid r}(X)\mu_{t}(X)\mathbb{I}_{\mathrm{ov}}(X)]$. Let $\mathcal{U}$ denote the uncertainty set including any custom constraints, e.g., $\mathcal{U}=\mathcal{U}_{bnd}\cap\mathcal{U}_{\text{lip}}$. We define the pessimistic robust lower bound $\underline{V}_{\mathrm{rob}}(\pi)\coloneqq V_{\mathrm{ov}}(\pi)+\underline{V}_{\mathrm{nov}}(\pi)$, where
@@
$\textstyle\underline{V}_{\mathrm{nov}}(\pi)\coloneqq\min_{q_{1\mid r}(X,A)\in\mathcal{U}}\sum_{r\in\{0,1\}}\mathbb{E}\!\left[\pi_{r}(X,A)\bigl\{\mu_{0}(X,A)+q_{1\mid r}(X,A)\tau(X,A)\bigr\}\mathbb{I}_{\mathrm{nov}}(X)\right].$
Here $\underline{V}_{\mathrm{rob}}(\pi)$ is the pessimistic robust lower bound; the corresponding optimistic bound replaces the minimum by a maximum.
In the binary-outcome, constant-bound case, the worst-case no-overlap term admits a simple closed form based on the sign of the relevant conditional mean; see Section 12.3.1. We therefore state only the more general interval-bounded robust LP here. We consider the case of continuous-valued outcomes in the example setting of the simple treatment parity–constrained program of Equation 4. We first study simple uncertainty sets, such as intervals, to deduce insights about the robust policy, with a more general reformulation in the appendix. For interval bounds, it is convenient to summarize each interval by its midpoint $m_{r}$ and radius $\rho_{r}$. We now characterize the policy that maximizes the pessimistic robust lower bound $\underline{V}_{\mathrm{rob}}(\pi)$.
Proposition 5.5 (Overlap-robust linear program)
Suppose that $r,t\in\{0,1\}$ and $q_{1\mid r}(\cdot,u)\in\mathcal{U}_{\mathrm{bnd}}$ for all $r\in\{0,1\}$, $u\in\{a,b\}$, so that $\underline{B}_{r}(x,u)\leq q_{1\mid r}(x,u)\leq\overline{B}_{r}(x,u)$. Define midpoint and radius $m_{r}(x,u):=\frac{1}{2}({\underline{B}_{r}(x,u)+\overline{B}_{r}(x,u)}),\;\rho_{r}(x,u):=\frac{1}{2}({\overline{B}_{r}(x,u)-\underline{B}_{r}(x,u)})$, objective-at-midpoint $c_{1}(\pi):=\sum_{r\in\{0,1\}}\mathbb{E}\!\left[\tau(X,A)\pi_{r}(X,A)m_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\right]$, and disparity on the overlap region:
$\Delta_{\mathrm{ov}}^{T}(\pi):=\mathbb{E}[T(\pi)\mathbb{I}_{\mathrm{ov}}(X)\mid A=a]-\mathbb{E}[T(\pi)\mathbb{I}_{\mathrm{ov}}(X)\mid A=b].$
Then the overlap-robust policy optimization problem is:
$\max_{\pi}\;$ $V_{\mathrm{ov}}(\pi)+\mathbb{E}[\mu_{0}(X,A)\mathbb{I}_{\mathrm{nov}}(X)]+c_{1}(\pi)-\sum_{r\in\{0,1\}}\mathbb{E}\!\left[|\tau(X,A)|\,\pi_{r}(X,A)\,\rho_{r}(X,A)\,\mathbb{I}_{\mathrm{nov}}(X)\right]$
s.t. $\sum_{r\in\{0,1\}}\Big\{\mathbb{E}[\pi_{r}(X,A)\overline{B}_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\mid A=a]-\mathbb{E}[\pi_{r}(X,A)\underline{B}_{r}(X,A)\mathbb{I}_{\mathrm{nov}}(X)\mid A=b]\Big\}+\Delta_{\mathrm{ov}}^{T}(\pi)\leq\epsilon.$
A key payoff of our modeling perspective is that the take-up / treatment-efficacy decomposition makes the overlap-robust problem less conservative: when recommendation overlap fails, we only need to robustly bound extrapolated treatment response to recommendation, rather than the downstream outcome model itself. Later in Section 6 we illustrate how this allows us to robustly improve current recommendations.
5.2 Extension: variance-sensitive constrained policy learning
We previously discussed policy optimization over unrestricted decision rules given estimates. We now turn to implementable policy learning over a restricted policy class under more general fairness constraints. Let $\Delta(\pi)\in\mathbb{R}^{K}$ denote a vector of identified disparity moments, and consider
$\max_{\pi\in\Pi}\{E[Y(\pi)]:E[\Delta(\pi)]\leq d\}.$
This formulation also enables local Pareto improvement around a reference policy $\pi_{0}$, by treating allowable value loss relative to $\pi_{0}$ as an additional constraint. To obtain tighter out-of-sample feasibility, we solve these constrained problems with a two-stage procedure that localizes around a first-stage solution and calibrates variance-based slacks. We discuss our two-stage procedure building on a constrained optimization oracle, which we instantiate with (Agarwal et al. 2018), but other methods with statistical guarantees on the Lagrangian regret also suffice.
A constrained policy optimization oracle.
We describe a specific procedure for constrained policy optimization that we leverage to concretize our results.
Agarwal et al. (2018) develops a general saddle-point procedure that optimizes such fairness-constrained problems over a hypothesis class with two key algorithmic tools: convexifying the space of policies via randomization, and saddle-point optimization over the convexified constrained problem. Such hypothesis classes include probabilistic classifiers that are good parametrizations of our causal encouragement policy.
They convexify $\Pi$ via randomized policies $Q\in\mathcal{P}(\Pi)$, where $\mathcal{P}(\Pi)$ is the set of distributions over $\Pi,$ i.e., a randomized classifier that samples a policy $\pi\sim Q$. Hence, they solve:
$\max_{Q\in\mathcal{P}(\Pi)}\min_{\lambda}\mathcal{L}(Q,\lambda),\text{ where }\mathcal{L}(Q,\lambda)\coloneqq\mathbb{E}[V(Q)]-\lambda^{\top}(\mathbb{E}[{\Delta}(Q)]-{d})$
Of course, the stochastic objective and constraints need to be estimated from data. We let $\hat{V},\hat{\delta}$ denote such estimates of the policy value and constraint scores.
We also require a feasibility margin $\epsilon_{k}$ that depends on concentration of the estimated constraints, so the sampled constraint vector is $\hat{d}_{k}=d_{k}-\epsilon_{k},$ for all $k$, ensuring that the constraint holds at nominal level $d_{k}$ out of sample.
We introduce some additional notation. Let $O=(X,A,R,T,Y)\sim P$ denote a fresh observationFor a deterministic policy $\pi$, let $v_{DR}(\pi):=v_{DR}(O;\pi,\eta_{0})$ denote the random doubly robust value contribution, and let $\Delta(\pi):=\Delta(O;\pi)\in\mathbb{R}^{K}$ denote the random vector of identified constraint contributions. Let $\delta(\pi):=\delta(O;\pi,\eta)$ denote the corresponding observed-data estimating score. We overload notation and discuss randomized policies via a distribution over policies $Q$, e.g. $v_{DR}(Q):=\mathbb{E}_{\pi\sim Q}\!\left[v_{DR}(\pi)\mid O\right],$ where $\mathbb{E}_{\pi\sim Q}[\cdot\mid O]$ denotes expectation over the policy randomization only, and use analogous variants. We seek an approximate saddle point so that the constrained solution is equivalent to the Lagrangian,
@@ 5.2 Extension: variance-sensitive constrained policy learning
$\textstyle\underset{Q\in\mathcal{P}(\Pi)}{\max}\;\;\underset{{{\lambda}\in\mathbb{R}_{+}^{K},\|{\lambda}\|_{1}\leq B}}{\min}\widehat{\mathcal{L}}(Q,{\lambda}),\qquad\underset{{\lambda}\in\mathbb{R}_{+}^{K},\|{\lambda}\|_{1}\leq B}{\min}\;\;\underset{Q\in\mathcal{P}(\Pi)}{\max}\;\widehat{\mathcal{L}}(Q,{\lambda})$
@@
If convexifying via randomization is not permitted, we discuss a simple generalization in the appendix that leverages well-known convex surrogate losses for optimizing optimal treatment regimes (Zhao et al. 2012). We defer specific details to Section 11.2.1, including how to encode fairness constraints in standard convexified LP form. [^1] The centered policy regret can be reparametrized via a policy parameter $\beta$ as: $J(\beta)=J(\operatorname{sgn}(f_{\beta}(\cdot)))=\mathbb{E}[\operatorname{sgn}(f_{\beta}(X))\left\{\psi\right\}].$ We can apply the standard reduction to cost-sensitive classification since $\psi_{i}\operatorname{sgn}(f_{\beta}(X_{i}))=\left|\psi_{i}\right|(1-2\mathbb{I}\left[\operatorname{sgn}(f_{\beta}(X_{i}))\neq\operatorname{sgn}(\psi_{i})\right]$). Then, we can use surrogate losses for the zero-one loss. Although many functional forms for $\ell(\cdot)$ are Fisher consistent, one such choice of $\ell$ is the logistic (cross-entropy) loss $\mathbb{E}[\left|\psi\right|\ell(f_{\beta}(X),\operatorname{sgn}(\psi))],\qquad l(g,s)=2\log(1+\exp(g))-(s+1)$.